OpenStudy (littleheron):

need help! will medal and fan! Solve for y. Show all your work -12x + 3y = 15

pooja195 (pooja195):

What do you think the first step would be?

OpenStudy (anonymous):

You want to get y isolated on one side of the equation.

OpenStudy (anonymous):

How would you do this?

OpenStudy (littleheron):

divide 3y from both sides?

pooja195 (pooja195):

Close... we do that AFTER we add 12x to both sides

OpenStudy (littleheron):

oh

pooja195 (pooja195):

\[\huge~\rm~\bf~-12x+12x + 3y = 15+12x\]

OpenStudy (anonymous):

@pooja195 is correct. Whatever operation is being done in an equation, we want to undo it. So if its being added we want to subtract, is it's being multiplied we want to divide, etc.

pooja195 (pooja195):

Now we have 3y=15+2x divide all the numbers by 3 what are you left with? (its gonna be an equation)

OpenStudy (littleheron):

1y = 5 + 1.5x?

pooja195 (pooja195):

Lets leave it in fraction form (the 3/2) But yes the answer is correct now to put it in slope intercept form y=mx+b we get y=2/3x+5 :) And that's all it is.

pooja195 (pooja195):

*2/3